AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 446     Entry time: Wed Sep 27 15:56:41 2023
Author: TD 
Subject: Wednesday 27 September 
16.55 CARME IE514 8.2e-11 mbar

      Vacuum pressure increased slowly overnight from c. 6e-11 mbar yesterday to c. 8e-11 mbar today following low power startup tests of the bakeout controller heating 
the CARME interaction and detector chambers yesterday and today.


Attachment 1 - bakeout controller config courtesy Anton Kalinin

Attachments 2 & 3 - bakeout controller summary displays and CARME bakeout monitor display c. 17.00 today

The 3x CARME heaters are connected to controller #4 zones # 1-3, i.e. lower right display, zones 1-3, mains fuses F9-F11

17.30 3x ( of 4x ) SAES D-2000 ion pump controllers ( SIP modules ) switched off. Fourth controller was already switched off and SHV cable disconnected from rear panel of 
SIP module.
Attachment 1: 1000007129.jpg  1.247 MB  Uploaded Wed Sep 27 17:05:17 2023  | Hide | Hide all | Show all
1000007129.jpg
Attachment 2: 1000007127.jpg  1.331 MB  Uploaded Wed Sep 27 17:05:17 2023  | Hide | Hide all | Show all
1000007127.jpg
Attachment 3: 1000007128.jpg  1.731 MB  Uploaded Wed Sep 27 17:05:17 2023  | Show | Hide all | Show all
ELOG V3.1.4-unknown